OpenStudy (anonymous):

Given that the two vectors 8u-v and 4u+3v are perpendicular, and the magnitude of v is 2u , find the angle bet?

OpenStudy (anonymous):

Use the fact that \[ \mathbf{a}\cdot \mathbf{b} = ||\mathbf{a}||\cdot ||\mathbf{b} ||\cdot \cos\theta \]The angle you want is \(\theta\).

OpenStudy (anonymous):

Wait a second....

OpenStudy (anonymous):

I know that, but they didn't give any number

OpenStudy (anonymous):

:[

OpenStudy (anonymous):

Use the fact that \[ (8\mathbf{u}-\mathbf{v}) \cdot (4\mathbf{u}+3\mathbf{v}) = 0 \]

OpenStudy (anonymous):

8u